Circular No. 7746

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2001 U1, C/2001 U2, C/2001 U3, C/2001 U4, C/2001 U5

Further to IAUC 7741, D. Hammer reports measurements of additional sungrazing comets visible in the C2 coronagraph and found on SOHO website images by X. Leprette (C/2001 U1, C/2001 U2), S. Hönig (C/2001 U3, C/2001 U4), and M. Boschat (C/2001 U5); C/2001 U3, C/2001 U4, and C/2001 U5 were also seen in C3 images. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 U1 Oct. 16.535  13^{h}18.^{m}0  − 9^{o}45′  2001-V10
C/2001 U2  17.368  13 21.1  −10 08  2001-V19
C/2001 U3  17.071  13 09.6  −10 40  2001-V19
C/2001 U4  17.013  13 01.4  −11 20  2001-V19
C/2001 U5  18.388  13 18.0  −11 05  2001-V19

COMET C/2001 U6 (LINEAR) An apparently asteroidal object discovered on Oct. 29 by LINEAR,

posted on the NEO Confirmation Page due to its unusual motion, has been found to be cometary by J. Tichá and M. Tichý (Kleť, diffuse with “a slight coma” on Nov. 3.8 UT) and by R. Trentman (Louisburg, KS, Nov. 6.3).

2001 UT  α_{2000}  δ_{2000}  m_{2}
Oct. 29.39508  8^{h}54^{m}53.^{s}11  +66^{o}34′22′′.2  19.5

The available astrometry, together with the following parabolic orbital elements and an ephemeris, appear on MPEC 2001-V27.

T = 2002 Aug. 19.637 TT  ω = 87.^{o}896
Ω = 115.383 }2000.0
q = 4.33741 AU  i = 107.423

SUPERNOVA 2001fu IN MCG –03-23-11 W. D. Li reports the discovery by LOTOSS (IAUC 7514) of an appar-

ent supernova (mag ∼ 15.1) in unfiltered images taken with the Katzman Automatic Imaging Telescope on Nov. 5.5 and 6.5 UT. SN 2001fu is located at α = 8^{h}52^{m}16^{s}.58, δ = −17^{o}44′29′′.8 (equinox 2000.0), which is 25′′.1 west and 10′′.6 north of the nucleus of MCG –03-23-11. A KAIT image taken on Apr. 15.3 showed nothing at this position (limiting mag ∼ 19.0).

2001 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green